A rectangular fence is to be enclosed by 400m of fence. What dimension will give a maximum area?
Answer by checkley77(12844)

THE MAXIMUM AREA FOR A RECTANGLE IS:
X^2=400
X=SQRT400
X=20 ANS. FOR THE SIDES OF THE SQUARE.
PROOF:
20*20=400
TEST OTHER MEASUREMENTS.
GIVEN X+Y=40.
X*Y<400
19*21=399
18*22=396
30*10=300